OpenStudy (brebre123):

Use substitution to solve the system. x = +2y8 −2x5y = 19

OpenStudy (brebre123):

Use substitution to solve the system. x= 2y+8 2x-5y=19

OpenStudy (brebre123):

@3mar ?

OpenStudy (3mar):

Well, I am here.

OpenStudy (brebre123):

can u help me please?

OpenStudy (3mar):

of course

OpenStudy (3mar):

What did you get?

OpenStudy (brebre123):

idk how to do it...

OpenStudy (3mar):

it is simple In Sha' Allah It depends on that you replace one variable in one equation with its value from the other equation.

OpenStudy (brebre123):

is it 2,-3?

OpenStudy (3mar):

ex: x+y=5 >>>>>>x=(y+5) x+2y=-1 >>replace x and put its value (y+5) y+5+2y=-1 then simplify etc

OpenStudy (3mar):

Are they x and y?

OpenStudy (3mar):

@brebre123